OpenStudy (ttp):

For which of the following systems of linear equations is the ordered pair negative one comma seven a solution? x minus four y equals one…two x minus three y equals eleven x plus two y equals five…x plus three y equals four x minus five y equals negative two…three x plus nine y equals twelve y equals eight plus x…x plus two y equals thirteen

OpenStudy (texaschic101):

we need to sub (-1,7) into your answer choices and see which is correct x - 4y = 1 -1 - 4(7) = 1 -1 - 28 = -1 -29 = -1 (not this one) we do not need to try the other equation because they both have to be true. x + 2y = 5 -1 + 2(7) = 5 -1 + 14 = 5 13 = 5 (not this one) x - 5y = -2 -1 - 5(7) = -2 -1 - 35 = -2 -36 = -2 (nope) y = 8 + x 7 = 8 + (-1) 7 = 8 - 1 7 = 7 (correct) now we will try the other equation x + 2y = 13 -1 + 2(7) = 13 -1 + 14 = 13 13 = 13 (correct) It is the last answer choice because (-1,7) satisfies both equations

OpenStudy (ttp):

Ok, :) I did one just similar to this could you check my answer ?

OpenStudy (texaschic101):

of course

OpenStudy (ttp):

A first grade class has 6-year olds and 7-year olds. There are 13 children in the class, and the total of their ages is 83. Which system of linear equations represents the problem? seven x plus six y equals thirteen x plus y equals eighty-three six x plus seven y equals eighty-three x plus y equals thirteen x six x plus seven y equals thirteen x plus y equals eighty-three x seven x plus six y equals eighty-three x plus y equals thirteen I said it was "B" Am i right? iF I'm not it's opition D i think ?

OpenStudy (texaschic101):

lets try B 6x + 7y = 83 x + y = 13x ----not this one....there is not supposed to be an x after 13 lets try D 7x + 6y = 83 x + y = 13 (this is correct) answer is D.

OpenStudy (ttp):

Okay thank you ! :)

OpenStudy (texaschic101):

sure :)

OpenStudy (ttp):

I have one more i neeed you to check

OpenStudy (texaschic101):

lets see it

OpenStudy (ttp):

Lines A and B form a linear system. Line A has an undefined slope and an x-intercept of (6, 0). Line B also has an undefined slope and has an x-intercept of (-4, 0). How are these lines characterized? parallel intersecting coincidental not enough information I said B

OpenStudy (texaschic101):

when the slope of the line is undefined, it is a vertical line. Is the slope undefined or is it 0 ?

OpenStudy (ttp):

Undefined

OpenStudy (texaschic101):

I am going to say parallel...but not 100% sure

OpenStudy (ttp):

oK :)